OpenStudy (anonymous):

I need to simplify this expression as much as possible with no negative numbers radicals and no radicals in denominators. could you help me please? the square root of -20 times the square root of 5

OpenStudy (anonymous):

\[\large \sqrt {-20} \times \sqrt {5} = i\sqrt {4 \times 5} \times \sqrt {5} = 2i \sqrt {5} \times \sqrt {5}=2i \times 5 = 10i \]